Can anyone show me step by step how to solve these? :(

4A)     In the reaction, 2 H2O2 → 2 H2O + O2

a.      How many grams of water are produced from 16.0 moles of hydrogen peroxide?

b.     How many grams of water are produced from 50.0 g of H2O2?

5A)    In the reaction, 3NaOH (aq) + H3PO4 (aq) → Na3PO4 (aq) + 3HOH (l)

a.      Which reactant is the limiting reactant if there are 10.0 mol of NaOH and 10.0 mol of H3PO4?

b.     Find the theoretical yield of water in grams?

6A)     In the reaction, CO (g) + 2 H2 (g) →CH3OH (l)

a. Which reactant is the limiting reactant and theoretical yield if 150.0 g of CO and 50.0 g of H2 are mixed and allowed to react?

b. What is the percent yield if you actually made 155.0 grams
Answer #1

Q4.

a)

ratio is 2 mol of H2O2 = 2 mol of H2O

so, 1:1

then, 16 mol of H2O2 should produce 16 mol of H2O

now, 1 mol of H2O = 18 g

so 16 mol of water = 16*18 = 288 g of H2O

b)

first, change to mol

1 mol of H2O2 = 34 g/mol

mol of H2O2 = mass/MW = 50/34 = 1.4705 mol

ratio is 1:1 so 1.4705 mol of H2O2 =1.4705 mol of H2O

mass ofwater = mol*MW = 1.4705*18 = 26.469 g of H2O

Q5.

ratio is 3:1

so 10 mol of H3PO4 requires 10xx3 = 30 mol of NaOH, which we do NOT have, therefore NaOH will limit the reaction

c)

theoretical yield -- from total mass expected

10 mol of NaOH should produce 10 mol of H2O

so

mass of H2O = 10*18 = 180 g of H2O expected

Q6.

limiting reactant depends mainly on mol amount

mol of CO = mass/MW = 150/28 = 5.357

mol of H2 = mass/MW = 50/2 = 25 mol

clearly, there is excess of H2

limiting reactant is CO

b.

% yield = actual amount / theoretical * 100%

theoretical amount of methanol -- 1 mol of CO = 1 mol of CH3OH

5.357 mol of CO = 5.357 mol of methanol

mass of methanol expected = mol*MW = 5.357*32 = 171.424 g of methanol theoretically

% yield = (155)/171.424 *100 = 90.42 %